The equation of a line is 2x-2 =6. What is the slope of the line? Would you please explain it to me. This is how far I got:
2x-2=6
2x=8
2/2 = 8/2
x = 4
Thank you

Answer by stanbon(75887)

The equation of a line is 2x-2 =6. What is the slope of the line? Would you please explain it to me. This is how far I got:
2x-2=6
2x=8
2/2 = 8/2
x = 4
------
Right.
The graph would be a vertical line passing thru
all the points where x = 4.
---
The slope is "undefined".
Cheers,
Stan H.
---------------